Convert the repeating decimal to a/b form, where a,b are intergers and b is not equal to 0.
2.18

Answer by Alan3354(69443)

Convert the repeating decimal to a/b form, where a,b are intergers and b is not equal to 0.
2.18
---------
n = 2.181818...
----
100n = 218.181818...
-1n = - 2.181818...
----------------------- subtract
99n = 216
n = 216/99 = 24/11